Circular No. 5867

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 RP J. Luu, Stanford University; and D. Jewitt, University of Hawaii, re-

port their discovery of another very faint object with retrograde opposition motion of ∼ 3′′/hr. The object, which appears stellar in 0′′.8 seeing, was detected in CCD images obtained with the University of Hawaii’s 2.2-m telescope on Mauna Kea.

1993 UT  α_{2000}  δ_{2000}  R
Sept. 15.4963  0^{h}49^{m}01.^{s}88  +5^{o}13′37′′.6  24.5
15.5850  0 49 01.42  +5 13 34.8
16.4558  0 48 57.17  +5 13 07.0
16.4976  0 48 56.97  +5 13 05.8
17.4630  0 48 52.23  +5 12 35.1
17.4964  0 48 52.06  +5 12 33.8

The circular direct orbit computed for this object by the undersigned gives a distance intermediate between that of 1993 RO (cf. IAUC 5865) and the earlier Jewitt-Luu discoveries but closer to the former. The distance would again change by several AU if a much more eccentric or a retrograde solution were adopted.

Epoch = 1993 Sept. 10.00 TT  u = 180.^{o}57
Ω = 192.15 }2000.0
a  = 35.374 AU  i = 2.79
1993TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Sept. 10 0^{h}49.^{m}47 + 5^{o}16.^{′}5 34.467 35.375 154.^{o}0 0.^{o}7 25.1
20 0 48.66 + 5 11.2 34.408 35.375 163.9 0.5 25.0
30 0 47.80 + 5 05.7 34.379 35.375 174.0 0.2 25.0
Oct. 10 0 46.92 + 5 00.0 34.378 35.375 175.9 0.1 25.0
20 0 46.06 + 4 54.3 34.408 35.375 165.8 0.4 25.0

POSSIBLE NOVA IN LUPUS W. Liller, Viña del Mar, Chile, reports his discovery of a possible nova

of mag 8.0 (average of two photographs, Tech Pan film + orange filter) on Sept. 19.02 UT at α = 14^{h}28^{m}.4, δ = −50^{o}57′ (equinox 1950.0). His examination of some 50 films taken during 1991 Jan.-1993 May 4 reveals no star brighter than mag 11.5 at this position.

1993 September 20 Brian G. Marsden